OpenStudy (yoloshroom):

rewrite as a single logarithm. 2+log_4(a)+1/2log_4(b)-3log_4(c)-log_4(d) what do i do with the 2

OpenStudy (mathstudent55):

You could try writing 2 as a log base 4.

OpenStudy (yoloshroom):

Still a bit confused. I know the final answer... it's log_4(16a(b^1/2)/c^3d

OpenStudy (mathstudent55):

\(\log_4 x = 2\) What is x?

OpenStudy (yoloshroom):

ah so x=2^4 so x = 16?

OpenStudy (mathstudent55):

Correct. Now rewrite \(2\) as \(\log_4 16\) Then you will have an expression with all logs base 4. Then you can use the rules of logs to write a single logarithm.

OpenStudy (mathstudent55):

\(2+\log_4 a+\dfrac{1}{2}\log_4 b - 3 \log_4 c - \log_4 d\) \(=\log_4 16+\log_4 a+\dfrac{1}{2}\log_4 b - 3 \log_4 c - \log_4 d\)

OpenStudy (yoloshroom):

i see thank you :)

OpenStudy (mathstudent55):

You're welcome. Do you know how to continue?

OpenStudy (yoloshroom):

yes i do, thank you very much :) was confused with the "2"

OpenStudy (mathstudent55):

\(=\log_4 16+\log_4 a+\log_4 b^{\frac{1}{2}} - (\log_4 c^3 + \log_4 d)\) \(=\log_4 16ab^{\frac{1}{2}} - \log_4 c^3d\) \(=\log_4 \dfrac{16ab^{\frac{1}{2}}}{c^3d}\) BTW, the answer you know is correct is indeed correct.

OpenStudy (yoloshroom):

thank you!